# Recent advances in text embedding: A Comprehensive Review of Top-Performing Methods on the MTEB Benchmark

### v1

@@ 6.2 The universality analysis
The pursuit of developing a unified model to address a multitude of downstream tasks has been long-standing (?). Despite attempting to be general-purpose in previous models such as (?, ?, ?), studies indicate that these embedding models struggle to generalize across tasks and domains (?). In this section, we study whether the MTEB top performing text embeddings are becoming more universal due to the increasing number and improved quality of diverse text datasets across different tasks (?, ?), good quality synthetic data generated by LLMs (?, ?) as well as larger backbones such as LLMs.
@@
SimCSE (2021) (?) is selected as the baseline method as it is one of the cornerstone work in text embedding which is cited and adopted by most of the recent works. The improvements over different tasks of the top performing text embeddings compared to the baseline method SimCSE is shown in Table 4. Each text embedding model’s performance is divided by the baseline performance in the table: 1 means the model has the same performance as the baseline, larger than 1 value means the model improves the performance of the baseline, smaller than 1 value means the baseline outperforms the model. For the averaged metric, all the top performing text embeddings outperforms the baseline with a considerable gap (SimCSE is ranked 101th place). However, the improvements across different individual tasks are heavily imbalanced:
@@
  • Classification tasks: The logistic regression classifier, with a maximum of 100 iterations, is trained using the train set embeddings and its performance is evaluated on the test set (?). It can be seen from Table 4 that all of these top 25 best performing models are better than the baseline method SimCSE with varied improvements between 9% and 21%.
@@
  • Clustering tasks: A mini-batch k-means model is trained on the embedded texts, utilizing a batch size of 32 and setting k to match the total number of unique labels with the v-measure as the metric (?). All of the top performing text embedding models outperform the baseline by around 35%-57% increase over the baseline performance.
@@
  • Pair Classification (Pair-C): Duplicate or paraphrase pairs with binary labels are embedded and the average precision score based on cosine similarity on text embeddings is used as the main metric (?). The performance of all the top performing text embedding models is superior (with varied improvements between 15% and 20%) to the baseline in the Pair Classification task.
@@
  • Reranking tasks: Given a query and a list of relevant and irrelevant reference texts, cosine similarity is used to compare the embeddings and rank the references with MAP being the main metric (?). The Reranking tasks show an improved performance (between 22% and 28%) from all MTEB leading text embedding models compared to the baseline.
@@
  • Retrieval tasks: Given a corpus, queries and a mapping for each query to relevant documents from the corpus, cosine similarity scores on the embeddings between query and documents are used to rank documents for each query, with nDCG@10 being the main metric (?). The most considerable enhancement in the top-rated text embedding models of MTEB is observed in Retrieval tasks, with the majority of these models more than doubling the performance of baseline model.
@@
  • Semantic Textual Similarity (STS) tasks: Given sentence pairs labeled with continuous scores with higher numbers indicating more similar sentences, Spearman correlation based on cosine similarity between sentence pair embeddings is main metric (?). The increase in performance is moderate in STS tasks compared to other tasks for all top performing MTEB text embedding models, with the best performing model increasing 8.4% over the baseline performance.
@@
  • Summarization tasks: Given human-written and machine-generated summaries, cosine similarity between embeddings of machine summary and human summary is used to score the machine summaries with Spearman correlation being the main metric (?). Unlike other tasks, most of the top performing text embedding models are not able to outperform the baseline performance on summarization tasks.
@@
From the results in Table 4, it can be seen that compared to the baseline text embedding SimCSE published in 2021, most the top 25 best performing MTEB text embedding models (mostly published in 2023 and 2024) are not remarkably better than the baseline on all tasks, especially on Summarization tasks. All the top 25 text embedding models are notably better than the baseline model on Retrieval, Reranking, Clustering and Pair Classification tasks, especially on Retrieval task. The proposed methodologies appear to primarily impact the performance of retrieval tasks. However, it might be related to the training and fine-tuning datasets used by the top performing models and their similarity to MTEB benchmarks. Popular datasets used by the top performing models include StackExchange, Reddit, S2ORC, NLI (?), FEVER (?), NQ (?, ?), HotpotQA (?), Quora (?), MSMARCO, etc. These datasets are similar to MTEB benchmark datasets especially on Retrieval and Clustering tasks. Apart from datasets similarity, there are many efforts made by the state of the art embeddings to deal with the asymmetric tasks such as Retrieval, including generation of more synthetic asymmetric datasets as in (?, ?), instruction based embeddings as in (?, ?, ?, ?, ?, ?, ?), asymmetric formatting as in (?) and so on. Generally speaking, The results from Table 4 show that: the overall performance on MTEB benchmark are improved considerably by recent advances in universal text embeddings especially on Retrieval tasks while the performance on Summarization task sees no notable improvement compared to the baseline method.
@@
In terms of universality on languages, most of these models are trained on specific languages, typically English, and do not inherently accommodate multilingual data. This lack of language universality restricts their application in global, multilingual contexts. In the work of (?, ?), the authors use proprietary LLMs to generate synthetic data for a diverse range of text embedding tasks in 93 languages, covering hundreds of thousands of embedding tasks, which shows good performance on high-resource languages. However, for low-resource languages, there is still room for improvement as current open-source LLMs are not adequately pre-trained on them. In terms of the universality on text length, MTEB has Sentence to Sentence (S2S) tasks as well as Paragraph to Paragraph (P2P) tasks where the former only compare titles, while the latter include both title and content (?). For clustering tasks, Arxiv, Biorxiv, Medrxiv, Reddit and StackExchange have both S2S and P2P version, where S2S tasks have short texts with on average 57-115 chars and P2P tasks have long texts with on average 728-1981 chars. Most top performing text embeddings have better performances on P2P tasks on Arxiv, Biorxiv, Medrxiv, Reddit. However, on StackExchange data, most top performing text embeddings have much better performance on S2S tasks. This might be more related to the informativeness nature of datasets instead of to the text length. Better benchmark datasets design related to text length is needed. For example, comparing the clustering performance on long text data before and after different extends of summarization could be an option.
@@ 6.3 Model efficiency analysis
Model sizes: In previous studies (?), it was found that the performance strongly correlates with model size, which can be identified in Figure 4. For example, when the parameters of SGPT increases from 1.3B to 5.8B, the performance increases from 56.2% to 58.93%. Such kind of scaling behavior encourages many studies to scale model size up in order to provide state of the art results across different embedding tasks. Recently, there are more and more models focus on generating text embeddings from LLMs because it does not need the contrastive pre-training step used in existing state of the art text embedding models as LLMs are extensively pre-trained on web-scale data already (?, ?, ?, ?, ?, ?). However, LLMs are computationally expensive, resource-intensive, and difficult to deploy in real-world applications, particularly on devices with limited processing power. Moreover, the marginal gains in performance do not always justify the substantial increase in parameter size, complexity and resource requirements. Additionally, we can see that when GritLM 7B is scaled up to GritLM 7x8B, the overall performance on MTEB benchmark decreases across all tasks (note that Grit models are both embedding and generation models). The performances of 7B parameters models vary a lot from 57.59% (sgpt-bloom-7b1-msmarco) to 67.56% (SFR-Embedding-Mistral) as shown in Figure 4, while jina-embeddings-v2-small-en achieves better performance than sgpt-bloom-7b1-msmarco with only 33M parameters. Furthermore, the two 1.2B models voyage-lite-02-instruct and google-gecko.text-embedding-preview-0409 demonstrate comparable or superior performances to most 7B LLMs based models, which suggests that there is significant room for enhancement in the efficiency of numerous state of the art text embedding models.
@@
Embedding sizes: Deploying text embedding involves two steps: a constant forward pass to compute the embedding, and its use for downstream applications (?, ?). The computation costs for the second step rise with the embedding dimensionality, data size, and label space, which can exceed the feature computation cost for large scale systems (?, ?). In some RAG systems where documents are stored as text embedding vectors, the embedding dimension is also related to the storage and memory cost, especially for large scale RAG systems. The top-performing text embedding dimension sizes vary from 256 to 4096, while the largest embedding dimension reported in MTEB English benchmark is 12288 from text-similarity-davinci-001 and text-search-davinci-001. MRL (?) and 2dMSE (?) propose new loss functions to allow first $m$ dimensions of the embedding to be independently capable of being a general purpose text embedding too. Among the top performing text embedding models, Gecko (?) embeddings are the most compact with google-gecko.text-embedding-preview-0409 (768 dimensions) and google-gecko-256.text-embedding-preview-0409 (256 dimensions).
@@
Max token sizes: The max token size limits the length of the input text to be embedded. When the input length exceeds the max token size, the most straightforward solution is to truncate the input text to the maximum allowed length. However, this approach has the drawback of eliminating potentially relevant text. An alternative strategy involves partitioning the input text into smaller chunks, embedding each chunk separately, and then combine the embeddings of all chunks. Although this method preserves the entirety of the input text, it reduces the efficiency of the embedding model. The max token sizes of top performing text embedding models in MTEB English benchmark vary from 512 to 32768. For BERT like based models, their max token size is usually 512, while text embedding models based on Mistral-7B have the max token size of 32768. To be noted that different LLMs may have different max token sizes. For example, Llama 2 (?) with 7B, 13B, and 70B parameters have a max token size of 4096, which shows that the max token size is not necessarily correlated with model parameter size. Further more, the max token size can be extended in various ways for both LLMs (?) and BERT like models (?).
@@ 6.4 Limitations
The complexity of comparing different models arises due to variations in numerous factors such as training data, back-bone model, loss function, training strategy, negative-sampling strategy, embedding dimension, among others. It is challenging to establish a fair comparison due to these differences. Few papers analyze the similarity between their training, pre-training or fine-tuning data and the MTEB benchmark datasets which makes it unclear whether MTEB test datasets are in-domain, partially in-domain or out-of-domain for these text embedding models. Many studies claim that the dataset diversity is important to achieve the universal text embeddings (?, ?, ?, ?). However, the current literature lacks a metric to accurately measure this dataset diversity, further complicating the issue. This gap in the literature underscores the need for a more rigorous approach to assessing dataset diversity in future studies.
@@
Instruction refers to the task instruction, which specifies a description of the task that the embedding will be used for (as shown in Equation 12) in order to build universal text embedding models that can generalize across a large variety of tasks (?, ?, ?). Many studies have shown that adding instructions has a considerable impact on the performance. However, there are several limitations. Firstly, the effectiveness of the instruction is highly dependent on its quality and specificity. If the instruction is vague or ambiguous, the model may fail to embed the text properly, leading to poor performance on the task. Additionally, creating precise and comprehensive instructions for every possible task can be a labor-intensive and time-consuming process. Secondly, the model’s ability to interpret and follow the instructions is limited by its current understanding of language, which may not perfectly align with human understanding. This could lead to misinterpretations and errors. Furthermore, the incorporating instructions into text embeddings increases the input length which can be computationally intensive, particularly for large datasets and large models. Finally, few papers explain how instruction impacts the text embedding for symmetric and asymmetric tasks and helps improve the performance theoretically. How out-of-domain instructions impact the model performance is not clear neither.
@@
Massive Text Embedding Benchmark (MTEB) is the most popular and used benchmark for universal text embeddings. There are several already identified limitations of MTEB including: lacking long texts datasets (most test datasets MTEB have fewer than 500 chars), task imbalance (15 datasets on Retrieval task, 12 datasets on Classification task while only 1 dataset for Summarization task), limited multi-languages evaluation datasets and no programming language (code) datasets (?). Understanding syntax thoroughly is essential for a text embedding model to accurately determine the relationships between words, which aids in achieving a level of language comprehension that mirrors human cognitive processes (?). The capacity of text embedding models to generalize across various syntactic contexts is not sufficiently examined in the existing benchmark. Therefore, to evaluate the proficiency of text embedding models in understanding syntax, it would be beneficial to incorporate more datasets that focus on syntactic aspects. The variety of datasets can certainly be enhanced. For instance, out of the 11 datasets used for clustering in MTEB, six originate from scientific articles published on platforms like Arxiv, Biorxiv, and Medrxiv. It would be beneficial to include datasets from different fields like finance, business, arts, culture, health, travel, and more to broaden the scope.
@@
Distance metrics $d(\cdot,\cdot)$ in vector spaces must obey certain axioms or geometric constraints (?, ?) including:
@@
Cosine similarity is widely used in the literature and MTEB benchmark to measure similarity between text embeddings, which also obeys symmetry and an analogue of the triangle inequality (?). However, psychological representations of similarity do not always obey these constraints. The authors from (?, ?) show that some important aspects of human judgments of item similarity can not be captured by some of the geometric axioms of vector spaces. Researchers from (?, ?) demonstrate that human relational similarity judgments violate the geometric constraints of symmetry and the triangle inequality. A famous example in terms of violation of symmetry is that people judge North Korea to be more similar to China than the other way around (?). Furthermore, the authors from (?) conclude that cosine-similarity can yield arbitrary and meaningless similarities. Compared to the term of distance or kernel, dissimilarity and similarity are more general terms, which do not have the constraints to be a metric or positive semi-definite (?, ?). New (dis)similarity measures that aligns better with human judgments could be an interesting and important future directions.

### v2

@@ 6.2 The universality analysis
The pursuit of developing a unified model to address a multitude of downstream tasks has been long-standing [12]. Despite attempting to be general-purpose in previous models such as [31, 32, 33], studies indicate that these embedding models struggle to generalize across tasks and domains [34]. In this section, we study whether the MTEB top performing text embeddings are becoming more universal due to the increasing number and improved quality of diverse text datasets across different tasks [35, 36], good quality synthetic data generated by LLMs [34, 16] as well as larger backbones such as LLMs.
@@
SimCSE (2021) [57] is selected as the baseline method as it is one of the cornerstone work in text embedding which is cited and adopted by most of the recent works. The improvements over different tasks of the top performing text embeddings compared to the baseline method SimCSE is shown in Table 4. Each text embedding model’s performance is divided by the baseline performance in the table: 1 means the model has the same performance as the baseline, larger than 1 value means the model improves the performance of the baseline, smaller than 1 value means the baseline outperforms the model. For the averaged metric, all the top performing text embeddings outperforms the baseline with a considerable gap (SimCSE is ranked 101th place). However, the improvements across different individual tasks are heavily imbalanced:
@@
  • Classification tasks: The logistic regression classifier, with a maximum of 100 iterations, is trained using the train set embeddings and its performance is evaluated on the test set [37]. It can be seen from Table 4 that all of these top 25 best performing models are better than the baseline method SimCSE with varied improvements between 9% and 21%.
@@
  • Clustering tasks: A mini-batch k-means model is trained on the embedded texts, utilizing a batch size of 32 and setting k to match the total number of unique labels with the v-measure as the metric [37]. All of the top performing text embedding models outperform the baseline by around 35%-57% increase over the baseline performance.
@@
  • Pair Classification (Pair-C): Duplicate or paraphrase pairs with binary labels are embedded and the average precision score based on cosine similarity on text embeddings is used as the main metric [37]. The performance of all the top performing text embedding models is superior (with varied improvements between 15% and 20%) to the baseline in the Pair Classification task.
@@
  • Reranking tasks: Given a query and a list of relevant and irrelevant reference texts, cosine similarity is used to compare the embeddings and rank the references with MAP being the main metric [37]. The Reranking tasks show an improved performance (between 22% and 28%) from all MTEB leading text embedding models compared to the baseline.
@@
  • Retrieval tasks: Given a corpus, queries and a mapping for each query to relevant documents from the corpus, cosine similarity scores on the embeddings between query and documents are used to rank documents for each query, with nDCG@10 being the main metric [37]. The most considerable enhancement in the top-rated text embedding models of MTEB is observed in Retrieval tasks, with the majority of these models more than doubling the performance of baseline model.
@@
  • Semantic Textual Similarity (STS) tasks: Given sentence pairs labeled with continuous scores with higher numbers indicating more similar sentences, Spearman correlation based on cosine similarity between sentence pair embeddings is main metric [37]. The increase in performance is moderate in STS tasks compared to other tasks for all top performing MTEB text embedding models, with the best performing model increasing 8.4% over the baseline performance.
@@
  • Summarization tasks: Given human-written and machine-generated summaries, cosine similarity between embeddings of machine summary and human summary is used to score the machine summaries with Spearman correlation being the main metric [37]. Unlike other tasks, most of the top performing text embedding models are not able to outperform the baseline performance on summarization tasks.
@@
From the results in Table 4, it can be seen that compared to the baseline text embedding SimCSE published in 2021, most the top 25 best performing MTEB text embedding models (mostly published in 2023 and 2024) are not remarkably better than the baseline on all tasks, especially on Summarization tasks. All the top 25 text embedding models are notably better than the baseline model on Retrieval, Reranking, Clustering and Pair Classification tasks, especially on Retrieval task. The proposed methodologies appear to primarily impact the performance of retrieval tasks. However, it might be related to the training and fine-tuning datasets used by the top performing models and their similarity to MTEB benchmarks. Popular datasets used by the top performing models include StackExchange, Reddit, S2ORC, NLI [57], FEVER [75], NQ [70, 71], HotpotQA [72], Quora [76], MSMARCO, etc. These datasets are similar to MTEB benchmark datasets especially on Retrieval and Clustering tasks. Apart from datasets similarity, there are many efforts made by the state of the art embeddings to deal with the asymmetric tasks such as Retrieval, including generation of more synthetic asymmetric datasets as in [16, 38], instruction based embeddings as in [48, 38, 16, 39, 49, 34, 35], asymmetric formatting as in [34] and so on. Generally speaking, The results from Table 4 show that: the overall performance on MTEB benchmark are improved considerably by recent advances in universal text embeddings especially on Retrieval tasks while the performance on Summarization task sees no notable improvement compared to the baseline method.
@@
In terms of universality on languages, most of these models are trained on specific languages, typically English, and do not inherently accommodate multilingual data. This lack of language universality restricts their application in global, multilingual contexts. In the work of [38, 16], the authors use proprietary LLMs to generate synthetic data for a diverse range of text embedding tasks in 93 languages, covering hundreds of thousands of embedding tasks, which shows good performance on high-resource languages. However, for low-resource languages, there is still room for improvement as current open-source LLMs are not adequately pre-trained on them. In terms of the universality on text length, MTEB has Sentence to Sentence (S2S) tasks as well as Paragraph to Paragraph (P2P) tasks where the former only compare titles, while the latter include both title and content [37]. For clustering tasks, Arxiv, Biorxiv, Medrxiv, Reddit and StackExchange have both S2S and P2P version, where S2S tasks have short texts with on average 57-115 chars and P2P tasks have long texts with on average 728-1981 chars. Most top performing text embeddings have better performances on P2P tasks on Arxiv, Biorxiv, Medrxiv, Reddit. However, on StackExchange data, most top performing text embeddings have much better performance on S2S tasks. This might be more related to the informativeness nature of datasets instead of to the text length. Better benchmark datasets design related to text length is needed. For example, comparing the clustering performance on long text data before and after different extends of summarization could be an option.
@@ 6.3 Model efficiency analysis
Model sizes: In previous studies [37], it was found that the performance strongly correlates with model size, which can be identified in Figure 4. For example, when the parameters of SGPT increases from 1.3B to 5.8B, the performance increases from 56.2% to 58.93%. Such kind of scaling behavior encourages many studies to scale model size up in order to provide state of the art results across different embedding tasks. Recently, there are more and more models focus on generating text embeddings from LLMs because it does not need the contrastive pre-training step used in existing state of the art text embedding models as LLMs are extensively pre-trained on web-scale data already [38, 49, 39, 51, 48, 34]. However, LLMs are computationally expensive, resource-intensive, and difficult to deploy in real-world applications, particularly on devices with limited processing power. Moreover, the marginal gains in performance do not always justify the substantial increase in parameter size, complexity and resource requirements. Additionally, we can see that when GritLM 7B is scaled up to GritLM 7x8B, the overall performance on MTEB benchmark decreases across all tasks (note that Grit models are both embedding and generation models). The performances of 7B parameters models vary a lot from 57.59% (sgpt-bloom-7b1-msmarco) to 67.56% (SFR-Embedding-Mistral) as shown in Figure 4, while jina-embeddings-v2-small-en achieves better performance than sgpt-bloom-7b1-msmarco with only 33M parameters. Furthermore, the two 1.2B models voyage-lite-02-instruct and google-gecko.text-embedding-preview-0409 demonstrate comparable or superior performances to most 7B LLMs based models, which suggests that there is significant room for enhancement in the efficiency of numerous state of the art text embedding models.
@@
Embedding sizes: Deploying text embedding involves two steps: a constant forward pass to compute the embedding, and its use for downstream applications [85, 86]. The computation costs for the second step rise with the embedding dimensionality, data size, and label space, which can exceed the feature computation cost for large scale systems [87, 88]. In some RAG systems where documents are stored as text embedding vectors, the embedding dimension is also related to the storage and memory cost, especially for large scale RAG systems. The top-performing text embedding dimension sizes vary from 256 to 4096, while the largest embedding dimension reported in MTEB English benchmark is 12288 from text-similarity-davinci-001 and text-search-davinci-001. MRL [62] and 2dMSE [63] propose new loss functions to allow first $m$ dimensions of the embedding to be independently capable of being a general purpose text embedding too. Among the top performing text embedding models, Gecko [34] embeddings are the most compact with google-gecko.text-embedding-preview-0409 (768 dimensions) and google-gecko-256.text-embedding-preview-0409 (256 dimensions).
@@
Max token sizes: The max token size limits the length of the input text to be embedded. When the input length exceeds the max token size, the most straightforward solution is to truncate the input text to the maximum allowed length. However, this approach has the drawback of eliminating potentially relevant text. An alternative strategy involves partitioning the input text into smaller chunks, embedding each chunk separately, then combining the embeddings of all chunks. Although this method preserves the entirety of the input text, it reduces the efficiency of the embedding model. The max token sizes of top performing text embedding models in MTEB English benchmark vary from 512 to 32768. For BERT like based models, their max token size is usually 512, while text embedding models based on Mistral-7B have the max token size of 32768. To be noted that different LLMs may have different max token sizes. For example, Llama 2 [115] with 7B, 13B, and 70B parameters have a max token size of 4096. Further more, the max token size can be extended in various ways for both LLMs [116] and BERT like models [117]. As MTEB lacks datasets with larger length, it is not clear how Max token sizes may impact the performance of universal text embedding models.
@@ 6.4 Limitations
The complexity of comparing different models arises due to variations in numerous factors such as training data, back-bone model, loss function, training strategy, negative-sampling strategy, embedding dimension, among others. It is challenging to establish a fair comparison due to these differences. Few papers analyze the similarity between their training, pre-training or fine-tuning data and the MTEB benchmark datasets which makes it unclear whether MTEB test datasets are in-domain, partially in-domain or out-of-domain for these text embedding models. Many studies claim that the dataset diversity is important to achieve the universal text embeddings [35, 12, 38, 34]. However, the current literature lacks a metric to accurately measure this dataset diversity, further complicating the issue. This gap in the literature underscores the need for a more rigorous approach to assessing dataset diversity in future studies.
@@
Instruction refers to the task instruction, which specifies a description of the task that the embedding will be used for (as shown in Equation 12) in order to build universal text embedding models that can generalize across a large variety of tasks [48, 16, 36]. Many studies have shown that adding instructions has a considerable impact on the performance. However, there are several limitations. Firstly, the effectiveness of the instruction is highly dependent on its quality and specificity. If the instruction is vague or ambiguous, the model may fail to embed the text properly, leading to poor performance on the task. Additionally, creating precise and comprehensive instructions for every possible task can be a labor-intensive and time-consuming process. Secondly, the model’s ability to interpret and follow the instructions is limited by its current understanding of language, which may not perfectly align with human understanding. This could lead to misinterpretations and errors. Furthermore, the incorporating instructions into text embeddings increases the input length which can be computationally intensive, particularly for large datasets and large models. Finally, few papers explain how instruction impacts the text embedding for symmetric and asymmetric tasks and helps improve the performance theoretically. How out-of-domain instructions impact the model performance is not clear neither.
@@
Massive Text Embedding Benchmark (MTEB) is the most popular and used benchmark for universal text embeddings. There are several already identified limitations of MTEB including: lacking long texts datasets (most test datasets MTEB have fewer than 500 chars), task imbalance (15 datasets on Retrieval task, 12 datasets on Classification task while only 1 dataset for Summarization task), limited multi-languages evaluation datasets and no programming language (code) datasets [37]. Understanding syntax thoroughly is essential for a text embedding model to accurately determine the relationships between words, which aids in achieving a level of language comprehension that mirrors human cognitive processes [118]. The capacity of text embedding models to generalize across various syntactic contexts is not sufficiently examined in the existing benchmark. Therefore, to evaluate the proficiency of text embedding models in understanding syntax, it would be beneficial to incorporate more datasets that focus on syntactic aspects. The variety of datasets can certainly be enhanced. For instance, out of the 11 datasets used for clustering in MTEB, six originate from scientific articles published on platforms like Arxiv, Biorxiv, and Medrxiv. It would be beneficial to include datasets from different fields like finance, business, arts, culture, health, travel, and more to broaden the scope.
@@
Distance metrics $d(\cdot,\cdot)$ in vector spaces must obey certain axioms or geometric constraints [119, 120] including:
@@
Cosine similarity is widely used in the literature and MTEB benchmark to measure similarity between text embeddings, which also obeys symmetry and an analogue of the triangle inequality [121]. However, psychological representations of similarity do not always obey these constraints. The authors from [122, 123] show that some important aspects of human judgments of item similarity can not be captured by some of the geometric axioms of vector spaces. Researchers from [124, 122] demonstrate that human relational similarity judgments violate the geometric constraints of symmetry and the triangle inequality. A famous example in terms of violation of symmetry is that people judge North Korea to be more similar to China than the other way around [124]. Furthermore, the authors from [10] conclude that cosine-similarity can yield arbitrary and meaningless similarities. Compared to the term of distance or kernel, dissimilarity and similarity are more general terms, which do not have the constraints to be a metric or positive semi-definite [125, 126]. New (dis)similarity measures that aligns better with human judgments could be an interesting and important future directions.